A house which is valued at 2,600, 000 appreciated at rate of 12% per year



I. What will be its value after two and half years?



ii. After how long will its value be 4200,000

1
Expert's answer
2021-12-05T22:49:51-0500

i)

"FV=PV(1+r)^n"

"FV=2600000(1+0.12)^{2.5}=3451583.66"


ii)

"n=\frac{ln(FV/PV)}{ln(1+r)}=\frac{ln(42/26)}{ln1.12}=4.23" years